Tom has recently started a toy business. If everything goes well, he hopes to make an operating

income of 20,000. Tom makes and sells only one toy truck. He plans to sell each toy truck for

$10 dollars. Tom has a variable cost of 5 dollars a toy and fixed costs of 25,000. He is new to his

business, so he is looking at different ways to map his costs. Tom has reached out to you for

consulting. As a consultant firm write tom a memo letting him know his breakeven point,

quantity of toys needed to be sold for desired operating income and describe ABC costing. Let

him know if he should invest in such a system with his current business needs. If not, should he

consider using such a system in the future?Two pages memo length (1000 words).Provide explanation of CVP analysis, breakeven point, activity-based costing with

respective recommendations to Tom. Include a self-made graph of CVP analysis.
Answer #1

MEMO

TO          :               Mr. Tom

FROM   :               Mr. XYZ

DATE:                    XX/XX/XXXX

SUBJECT:             Project feasibility using CVP analysis ( Breakeven point & Activity based Costing)

We have studied and analyzed the data provided by you about the new business of toys and as per your requirements we have prepared a detailed summary of the project’s breakeven point (BEP), targeted sales quantity & explanation in respect to the Activity based costing.

What CVP Analysis is:

CVP analysis involves the study of various interrelated factors such as revenues, costs, levels of activity, profits and other factors. It is primarily used for planning & decision making but it can also be used for controlling and evaluating decisions.

By evaluating the information given by you such as Variable cost, Sales price, desired profit, etc. we have done a CVP analysis to evaluate the project and give you exact figures for the information required.

Data provided to us:

Desired profit    = $20000

Variable cost      = $5 per unit

Sales price           = $10 per unit

Fixed cost            = $25000

Calculation of Breakeven point

Break even point= Fixed Cost/contribution

Where, contribution= Sales price- Variable Cost

                                        = 10-5 = $5

Break Even point= 20000/5= 5000 Units of toys OR $50000

Quantity of sales for Desired profit

Sales required= (Fixed Cost + Desired profit)/Contribution

Sales Required= (25000+20000)/5

                                = 9000 Units of Toys

Activity Based Costing

In activity based costing, costs are divided into categories of Unit-based and non-unit based costing. Most of the non-units based activities are fixed but many may vary when identifiable with a separate cost driver. Activity based CVP is a better option than conventional CVP as it takes into account the Variable Fixed cost or Incremental ficed cost which changes with the level of activity.

For Example,

Setup Cost increases when a new setup is used but it is done after a certain level of activity.

Activity Based Costing production of toys

The breakeven point is based on the fixed cost but with an increment in the level of production, the fixed cost might increase and thus there would be a new break even point. Say, new machinery is added which is a type of fixed cost.

Activity based costing helps in establishing the relation between data and the level of activity thus providing more accurate results comparatively.

Suggestion for the business:

The given information and the analysis done by us gives a positive result and favorable expectations from the business. Thus, it would be appropriate to go for the said project.

desped sales Area of Profit $90,000 Milli Total cost Breakeven ! Revenue a point sales Revenue a Area of Loss - 9000 5000 out